If the measure of an angle is twenty less than the measure of its complement, what is the measure of the angle?
Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

complement=x
angle=x-20
They add to 90
x+x-20=90
2x=110
x=55
x-20=35 degrees, the angle.

Answer by ikleyn(52787)

.
Let  x  be the measure of the given angle in degrees.
Then the measure of its complement is  x+20  degrees.

The sum of an angle and its complement is always 90°.

So you have an equation

x + (x + 20) = 90°.

2x = 90 - 20 = 70°.

x = 35°.

Answer.  35°.